# Proximality and selflessness for group C*-algebras

### v1

@@ 3. A space with “tree structure”
Let $\tilde{\Gamma}\coloneq\Gamma*\mathopen{\langle}z\mathclose{\rangle}$. By the normal form of $g\in\tilde{\Gamma}$, we mean the expression
$g=a_{1}z^{\epsilon_{1}}a_{2}z^{\epsilon_{2}}\cdots z^{\epsilon_{l}}a_{l+1}$
with $a_{i}\in\Gamma$ and $\epsilon_{i}\in\{\pm\}$, which is reduced in the sense that if $i>1$ and $a_{i}=1$, then $\epsilon_{i-1}=\epsilon_{i}$.
@@
We introduce the tree $\mathrm{T}\coloneq\tilde{\Gamma}/\Gamma$ on which $\tilde{\Gamma}$ acts. We declare that the points $s,t\in\mathrm{T}$ are adjacent if $s^{-1}t\in\{\Gamma z^{\pm}\Gamma\}$. With this graph structure $\mathrm{T}$ becomes a tree (of infinite degree). We call the point $o\coloneq\Gamma\in\mathrm{T}$ the origin. Every $t\in\mathrm{T}$ has unique representative $\rho(t)\in\tilde{\Gamma}$ whose normal form ends by $z^{\pm}$, or $\rho(o)=1$ if $t=o$. Note that $\rho(at)=a\rho(t)$ for all $a\in\Gamma$ and $t\in\mathrm{T}\setminus\{o\}$ and that $\rho(z^{\pm}t)=z^{\pm}\rho(t)$ for all $t\in\mathrm{T}$.
@@
Suppose first that $s$ and $t$ are adjacent. We label the oriented edge from $t$ to $s$ by $\ell(t,s)\coloneq\kappa(\rho(t)^{-1}s)\in\Gamma\{z_{\pm}\}$, where $\kappa(az^{\pm}\Gamma)\coloneq az_{\pm}$. Note that $t$ and $\ell(t,s)$ uniquely determines $s$. E.g., $\ell(az\Gamma,azbz\Gamma)=bz_{+}$ and $\ell(azbz\Gamma,az\Gamma)=z_{-}$ for every $a,b\in\Gamma$. The Figure 1 shows how $\mathrm{T}$ and $\ell$ look like. Next we extend the label on $(\mathrm{T}\times\bar{\mathrm{T}})\setminus\{\mbox{diagonal}\}$ by setting $\ell(t,s)\coloneq\ell(t,t_{1})$, where $t_{1}$ is the unique point between $t$ and $s$ that is adjacent to $t$, or colloquially, $\ell(t,s)$ is the first $\Gamma\{z^{\pm}\}$ component in the normal form $\rho(t)^{-1}\rho(s)$. For example, $\ell(az,azbzcz\Gamma)=bz_{+}$ and $\ell(cz\Gamma,azbz\Gamma)=z_{-}$ unless $a=c$.
@@
Let $(t,s)\in(\mathrm{T}\times\bar{\mathrm{T}})\setminus\{\mbox{diagonal}\}$. For $a\in\Gamma$, one has $\ell(at,as)=\ell(t,s)$ for $t\neq o$ and $\ell(o,as)=a\ell(o,s)$. Also, $\ell(zt,zs)=\ell(t,s)$.
Let $\Gamma\curvearrowright X$ be a topologically-free extreme boundary with an axial sequence $(z_{n})_{n}$. We denote by $\varphi_{n}\colon\tilde{\Gamma}\to\Gamma$ the homomorphism that is identity on $\Gamma$ and send $z$ to $z_{n}\in\Gamma$.
@@
Let $g=a_{1}z^{\epsilon_{1}}a_{2}z^{\epsilon_{2}}\cdots z^{\epsilon_{l}}$ be an element in the normal form with $l\geq 1$ and $x\in X\setminus\{z_{\bar{\epsilon}_{l}}\}$. Then $\varphi_{n}(g)x\to a_{1}z_{\epsilon_{1}}$ as $n\to\infty$.
Proof.
Let $U_{\pm}$ be arbitrarily small neighborhoods of $z_{\pm}$ that satisfy $x\notin U_{\bar{\epsilon}_{l}}$ and $a_{i}(U_{+}\cup U_{-})\cap(U_{+}\cup U_{-})=\emptyset$. Then by the ping-pong argument, one has $\varphi_{n}(g)x\in a_{1}U_{\epsilon_{1}}$ for all $n$ large enough. ∎
We denote by $\sigma$ the corresponding action of $\Gamma$ on $C(X)\subset C(X)^{\mathcal{U}}$, where the embedding is the constant embedding. We denote by $\sigma_{z}$ the automorphism arising from the sequence $(\sigma_{z_{n}})_{n}$. They together give rise to the action of $\tilde{\Gamma}=\Gamma*\mathopen{\langle}z\mathclose{\rangle}$ on $C(X)^{\mathcal{U}}$, which is given by
$\sigma_{g}([f_{n}]_{n})=[\sigma_{\varphi_{n}(g)}(f_{n})]_{n}.$
We are interested in the commutative $\mathrm{C}^{*}$-algebra
$C(X_{\mathrm{T}})\coloneq\mathrm{C}^{*}(\bigcup_{g\in\tilde{\Gamma}}\sigma_{g}(C(X)))\subset C(X)^{\mathcal{U}}$
on which $\tilde{\Gamma}$ acts, where $X_{\mathrm{T}}$ is the Gelfand spectrum of $C(X_{\mathrm{T}})$. By definition, there is a surjective homomorphism
@@
where $f^{(s)}\in C(X^{\mathrm{T}})$ is defined for $f\in C(X)$ and $\mathbf{x}=(x_{t})_{t}$ by $f^{(s)}(\mathbf{x})\coloneq f(x_{s})$. This gives rise to an embedding $Q_{*}\colon X_{T}\to X^{\mathrm{T}}$. We equip $X^{\mathrm{T}}$ a $\tilde{\Gamma}-$action by declaring that $a\in\Gamma$ acts at $\mathbf{x}=(x_{t})_{t}\in X^{\mathrm{T}}$ by $(a\mathbf{x})_{o}\coloneq ax_{o}$ and $(a\mathbf{x})_{t}\coloneq x_{a^{-1}t}$ for $t\in\mathrm{T}\setminus\{o\}$, and that $z$ acts by $(z\mathbf{x})_{t}\coloneq x_{z^{-1}t}$.
@@
Hereafter, we omit writing $Q$ and identify $\varsigma$ with $\sigma$. Thus we view $X_{\mathrm{T}}$ as a compact subset of $X^{\mathrm{T}}$. For every $s\in\mathrm{T}$, we define the continuous embedding $\theta(s,\,\cdot\,)\colon X\to X^{\mathrm{T}}$ by $\theta(s,x)_{s}\coloneq x$ and $\theta(s,x)_{t}\coloneq\ell(t,s)$ for $t\neq s$. We also define the continuous embedding $\theta\colon\partial\mathrm{T}\to X^{\mathrm{T}}$ by $\theta(\omega)_{t}\coloneq\ell(t,\omega)$. Note that $\omega\to\theta(\omega)_{t}$ is locally constant.
The following theorem asserts that $\Gamma\curvearrowright X_{\mathrm{T}}$ has a “tree structure”. See Figure 2.
Figure 2. The space $X_{\mathrm{T}}$ around $\theta(s,X)$, where $s=az\Gamma$ and $t=azbz\Gamma$, $\theta(s,bz_{+})=\theta(t,z_{-})$ and $\theta(s,z_{-})=\theta(o,az)$.
@@
In the above setting for $\tilde{\Gamma}=\Gamma*\mathopen{\langle}z\mathclose{\rangle}$, the following hold.
@@
The “if” part of (7-8) are easy. For the “only if” part, suppose that $(\theta(t_{n},x_{n}))_{n}$ converges to $\theta(s,x)$. If $(t_{n})_{n}$ is converges to some point $s^{\prime}$, then by the “if” part and the compactness of $X$, one has $\theta(t_{n},x_{n})\to\theta(s^{\prime},y)$, where $y$ is the limit of $y_{n}\coloneq x_{n}$ if $t_{n}=s^{\prime}$ and $y_{n}\coloneq\ell(s^{\prime},t_{n})$ if $t_{n}\neq s^{\prime}$. Since $\theta(s^{\prime},y)=\theta(s,x)$, one has either $s^{\prime}=s$ or $s^{\prime}$ is an adjacent point to $s$ that satisfies $x=\ell(s,s^{\prime})$ and $y=\ell(s^{\prime},s)$. This proves the “only if” part for the case $(t_{n})_{n}$ is convergent. Now consider the case where $(t_{n})_{n}$ is not convergent. By the previous case, the only limit points are either $s$ or the unique adjacent point $s^{\prime}$ that satisfies $x=\ell(s,s^{\prime})$. The other claimed conditions also follow. The rest of the proof of (7-8) is similar.
@@
It remains to prove the opposite inclusion. Let $\mathbf{x}\coloneq(x_{t})_{t}\in X^{\mathrm{T}}$ be given and assume that $\mathbf{x}\in X_{\mathrm{T}}$. We define $w\colon\mathrm{T}\to\mathrm{T}$, “the orientation of $\mathbf{x}$ at tt”, as follows. If $x_{t}\in X\setminus\Gamma\{z_{\pm}\}$, then $w(t)\coloneq t$. If $x_{t}\in\Gamma\{z_{\pm}\}$, then $w(t)\coloneq\rho(t)\kappa^{-1}(x_{t})$, where $\kappa^{-1}(az_{\pm})\coloneq az^{\pm}\Gamma\in\mathrm{T}$. Note that $w(t)$ is adjacent to $t$ in the latter case. We claim that if $s,t\in\mathrm{T}$ are adjacent, then it cannot happen that $w(s)\neq t$ and $w(t)\neq s$ simultaneously. Suppose for a contradiction that it happened. We may assume that $s$ is between $o$ and $t$ and $\rho(t)=\rho(s)az$ with no cancellation. Thus one has $x_{s}\neq az_{+}$ and $x_{t}\neq z_{-}$. Let $f\in C(X)$ be such that $f(x_{s})=1$ and $f=0$ on a neighborhood of $az_{+}$. Similarly let $g$ be such that $g(x_{t})=1$ and $g=0$ on a neighborhood of $z_{-}$. We consider the function $f^{(s)}g^{(t)}$ on $X^{\mathrm{T}}$. It satisfies $(f^{(s)}g^{(t)})(\mathbf{x})=f(x_{s})g(x_{t})=1$. However,

### v6

@@ 3. Extreme boundaries and “tree-graded” spaces
Let $\tilde{\Gamma}\coloneq\Gamma*\mathopen{\langle}z\mathclose{\rangle}$. By the reduced form of $g\in\tilde{\Gamma}$, we mean the unique expression
$g=a_{1}z^{\epsilon_{1}}a_{2}z^{\epsilon_{2}}\cdots a_{l}z^{\epsilon_{l}}a_{l+1}$
@@
with $a_{i}\in\Gamma$ and $\epsilon_{i}\in\{\pm\}$ such that $i>1$ and $a_{i}=1$ implies $\epsilon_{i-1}=\epsilon_{i}$. We introduce the tree structure on $\mathrm{T}\coloneq\tilde{\Gamma}/\Gamma$ on which $\tilde{\Gamma}$ acts. We declare that the points $s,t\in\mathrm{T}$ are adjacent if $s^{-1}t\in\{\Gamma z^{\pm}\Gamma\}$. With this graph structure $\mathrm{T}$ becomes a tree (of infinite degree). We call the point $o\coloneq\Gamma\in\mathrm{T}$ the origin. Every $t\in\mathrm{T}$ has unique representative $\rho(t)\in\tilde{\Gamma}$ whose reduced form ends by $z^{\pm}$, or $\rho(o)=1$ if $t=o$. Note that $\rho(at)=a\rho(t)$ for all $a\in\Gamma$ and $t\in\mathrm{T}\setminus\{o\}$ and that $\rho(z^{\pm}t)=z^{\pm}\rho(t)$ for all $t\in\mathrm{T}$.
@@
Suppose first that $s$ and $t$ are adjacent. We label the oriented edge from $t$ to $s$ by $\ell(t,s)\coloneq\kappa(\rho(t)^{-1}s)\in\Gamma\{z^{\pm}\}$, where $\kappa(az^{\pm}\Gamma)\coloneq az^{\pm}$. Note that $t$ and $\ell(t,s)$ together uniquely determines $s$. E.g., $\ell(az\Gamma,azbz\Gamma)=bz$ and $\ell(azbz\Gamma,az\Gamma)=z^{-1}$ for every $a,b\in\Gamma$. The Figure 1 shows how $\mathrm{T}$ and $\ell$ look like. Next we extend the label map on $(\mathrm{T}\times\bar{\mathrm{T}})\setminus\{\mbox{diagonal}\}$ by setting $\ell(t,s)\coloneq\ell(t,t_{1})$, where $t_{1}$ is the unique point between $t$ and $s$ that is adjacent to $t$, or colloquially, $\ell(t,s)$ is the first $\Gamma\{z^{\pm}\}$ component in the reduced form of $\rho(t)^{-1}\rho(s)$. For example, $\ell(az,azbzcz\Gamma)=bz$ and $\ell(cz\Gamma,azbz\Gamma)=z^{-1}$ unless $a=c$.
@@
Let $(t,s)\in(\mathrm{T}\times\bar{\mathrm{T}})$ be a non-diagonal point. For $a\in\Gamma$, one has $\ell(o,as)=a\ell(o,s)$ and $\ell(at,as)=\ell(t,s)$ for $t\neq o$. Also, $\ell(zt,zs)=\ell(t,s)$.
Let $\Gamma\curvearrowright X$ be an extreme boundary with an axial sequence $(z_{n})_{n}$. We denote by $\varphi_{n}\colon\tilde{\Gamma}\to\Gamma$ the homomorphism that is identity on $\Gamma$ and sends $z$ to $z_{n}\in\Gamma$. A simple ping-pong argument on $X$ shows that $(\varphi_{n})_{n}$ is asymptotically injective and hence $\Gamma$ is mixed-identity-free. More precisely, the following is true.
@@
Let $g=a_{1}z^{\epsilon_{1}}\cdots z^{\epsilon_{l}}a_{l+1}$ be an element in the reduced form with $l\geq 1$ and $x\in X\setminus\{a_{l+1}^{-1}z_{\bar{\epsilon}_{l}}\}$, where $\bar{\epsilon}$ is the opposite of $\epsilon$. Then $\varphi_{n}(g)x\to a_{1}z_{\epsilon_{1}}$ as $n\to\infty$.
We denote by $\sigma$ the corresponding action of $\Gamma$ on $C(X)\subset C(X)^{\mathcal{U}}$, where the embedding is the diagonal embedding. We denote by $\sigma_{z}$ the automorphism arising from the sequence $(\sigma_{z_{n}})_{n}$. They together give rise to the action of $\tilde{\Gamma}=\Gamma*\mathopen{\langle}z\mathclose{\rangle}$ on $C(X)^{\mathcal{U}}$, which is given by $\sigma_{g}([f_{n}]_{n})=[\sigma_{\varphi_{n}(g)}(f_{n})]_{n}$. We are interested in the $\tilde{\Gamma}-$invariant $\mathrm{C}^{*}$-subalgebra
$C(X_{\mathrm{T}})\coloneq\mathrm{C}^{*}(\bigcup_{g\in\tilde{\Gamma}}\sigma_{g}(C(X)))\subset C(X)^{\mathcal{U}},$
where $X_{\mathrm{T}}$ is the Gelfand spectrum of $C(X_{\mathrm{T}})$. By definition, there is a surjective homomorphism
@@
where $f^{(s)}\in C(X^{\mathrm{T}})$ is defined for $f\in C(X)$ by $f^{(s)}(\mathbf{x})\coloneq f(x_{s})$, $\mathbf{x}=(x_{t})_{t}$. This gives rise to an embedding $Q_{*}\colon X_{T}\to X^{\mathrm{T}}$. We equip $X^{\mathrm{T}}$ a $\tilde{\Gamma}-$action by declaring that $a\in\Gamma$ acts at $\mathbf{x}=(x_{t})_{t}\in X^{\mathrm{T}}$ by $(a\mathbf{x})_{o}\coloneq ax_{o}$ and $(a\mathbf{x})_{t}\coloneq x_{a^{-1}t}$ for $t\in\mathrm{T}\setminus\{o\}$, and that $z$ acts by $(z\mathbf{x})_{t}\coloneq x_{z^{-1}t}$.
@@
Hereafter, we omit writing $Q$ and identify $\varsigma$ with $\sigma$. Thus we view $X_{\mathrm{T}}$ as a compact subset of $X^{\mathrm{T}}$. We view the label map $\ell$ as a map taking values in $X$ by identifying $\Gamma\{z^{\pm}\}$ with $\Gamma\{z_{\pm}\}$. For each $s\in\mathrm{T}$, we define the continuous embedding $\theta(s,\,\cdot\,)\colon X\to X^{\mathrm{T}}$ by $\theta(s,x)_{s}\coloneq x$ and $\theta(s,x)_{t}\coloneq\ell(t,s)$ for $t\neq s$. We also define the continuous embedding $\theta\colon\partial\mathrm{T}\to X^{\mathrm{T}}$ by $\theta(\omega)_{t}\coloneq\ell(t,\omega)$. Note that $\omega\to\theta(\omega)_{t}$ is locally constant for every $t$.
The space $X_{\mathrm{T}}$ has a kind of “tree-graded” structure ([DS]), as follows. See Figure 2.
Figure 2. The space $X_{\mathrm{T}}$ around $\theta(az\Gamma,bz_{+})=\theta(azbz\Gamma,z_{-})$.
@@
In the above setting for $\tilde{\Gamma}=\Gamma*\mathopen{\langle}z\mathclose{\rangle}\curvearrowright\mathrm{T}$, the following hold.
@@
The “if” part of (7-8) are easy. For the “only if” part, suppose that $(\theta(t_{n},x_{n}))_{n}$ converges to $\theta(s,x)$. If $(t_{n})_{n}$ is converges to some point $s^{\prime}$, then by the “if” part and the compactness of $X$, one has $\theta(t_{n},x_{n})\to\theta(s^{\prime},y)$, where $y$ is the limit of $y_{n}\coloneq x_{n}$ if $t_{n}=s^{\prime}$ and $y_{n}\coloneq\ell(s^{\prime},t_{n})$ if $t_{n}\neq s^{\prime}$. Since $\theta(s^{\prime},y)=\theta(s,x)$, one has by (3) either $s^{\prime}=s$ or $s^{\prime}$ is an adjacent point to $s$ that satisfies $x=\ell(s,s^{\prime})$ and $y=\ell(s^{\prime},s)$. This proves the “only if” part for the case $(t_{n})_{n}$ is convergent. Now consider the case where $(t_{n})_{n}$ is not convergent. By the previous case, the only limit points are either $s$ or the unique adjacent point $s^{\prime}$ that satisfies $x=\ell(s,s^{\prime})$. The other claimed conditions also follow. The rest of the proof of (7-8) is similar.
@@
It remains to prove the opposite inclusion. Let $\mathbf{x}\coloneq(x_{t})_{t}\in X^{\mathrm{T}}$ be given and assume that $\mathbf{x}\in X_{\mathrm{T}}$. We define $w\colon\mathrm{T}\to\mathrm{T}$, “the orientation of $\mathbf{x}$ at tt”, as follows. If $x_{t}\in X\setminus\Gamma\{z_{\pm}\}$, then $w(t)\coloneq t$. If $x_{t}\in\Gamma\{z_{\pm}\}$, then we define $w(t)$ to be the unique point that is adjacent to $t$ and satisfies $\ell(t,w(t))=x_{t}$. We claim that if $s,t\in\mathrm{T}$ are adjacent, then it cannot happen that $w(s)\neq t$ and $w(t)\neq s$ simultaneously. Suppose for a contradiction that it happened. We may assume that $s$ is between $o$ and $t$ and $\rho(t)=\rho(s)az$ with no cancellation. Thus one has $x_{s}\neq az_{+}$ and $x_{t}\neq z_{-}$. Let $f\in C(X)$ be such that $f(x_{s})=1$ and $f=0$ on a neighborhood of $az_{+}$. Similarly let $g$ be such that $g(x_{t})=1$ and $g=0$ on a neighborhood of $z_{-}$. We consider the function $f^{(s)}g^{(t)}$ on $X^{\mathrm{T}}$. It satisfies $(f^{(s)}g^{(t)})(\mathbf{x})=f(x_{s})g(x_{t})=1$. However,